OpenStudy (colcaps):

tan 2B = cot (B - 12)?

OpenStudy (colcaps):

Its either 26 52 or 38

OpenStudy (anonymous):

tan(a+b)=tana+tanb/1-tana*tanb

OpenStudy (anonymous):

\[\frac{ 1 }{ \tan \theta} =\cot \theta\]

OpenStudy (colcaps):

well i have choices in my home work 26 52 38 i cant figure it out i think i have to simplify it

OpenStudy (anonymous):

\[\tan \theta=\cot (90-\theta)\]

OpenStudy (colcaps):

26 52 38 are my only options do i have to substitute?

OpenStudy (anonymous):

\[ \tan 2B=\tan (90-B+12)\]

hartnn (hartnn):

cot(B-12)= tan (90-B-12) so 90-B-12 =2B

OpenStudy (anonymous):

\[2B=102-B\]

hartnn (hartnn):

3B=102 B=102/3 = ?

OpenStudy (colcaps):

34

OpenStudy (colcaps):

close enough but thanks